# thanks: A footnote to the article title

### v2

@@ IV Conclusions
Despite its excellent performance in other areas,22, 154 the LAK meta-GGA has been found less accurate than r2SCAN for the oxidation energies (Fig. 2) and only slightly more accurate than r2SCAN for the band gaps (Fig. 5) of these transition-metal oxides.
Note that the first, second, and third ionization energies of the 3d transition metal atoms also improve from LSDA to PBE to r2SCAN.174 That improvement is dramatic for the second and third ionization energies, which seem free from strong correlation effects.
The present investigation and other recent works82, 83, 84, 91, 93, 174 suggest that, in many but not all situations with main-group and transition-metal elements, the functional-driven errors of r2SCAN energy differences can far exceed the relatively small density-driven errors. This has the important consequence that significant improvements in the r2SCAN energy differences can be achieved by doing the expensive self-consistent iteration and geometry optimization with the computationally efficient r2SCAN meta-GGA (possibly including an efficient van der Waals dispersion corrections), and then applying a more expensive nonlocal functional (e.g., a hybrid or self-interaction correction) in a single post-self-consistent calculation, as we have done here in r2SCAN10@r2SCAN. All properties studied here improve on average from LSDA to PBE to r2SCAN to r2SCAN10@r2SCAN. Replacing a small fraction of r2SCAN exchange by the same fraction of exact Hartree-Fock exchange should have little or no effect on the satisfaction of exact constraints.
@@ V First-Principles Calculation Details
All calculations presented in this work are performed using the DFT formalism, as implemented in the Vienna Ab initio Simulation Package (VASP).175, 176, 177 The PAW potentials describe the core electrons.178, 179 The electrons from 3s, 3p, 3d, and 4s orbitals are explicitly considered for the transition metal atoms. Using a PAW set that treats fewer valence electrons explicitly (3s, 3p, and 3d) significantly impacts MiOj oxidation energies and band gaps (Supplementary Fig. 8). The kinetic energy cutoff for the plane waves was set to 700 eV, and the total energy was converged to 10-6 eV per cell. Various DFT exchange and correlation functionals were used in their collinear spin-polarized implementation. These are the LSDA-PW92,17 the Perdew, Burke, and Ernzerhof (PBE),18 the meta-GGA functionals r2SCAN,99 and LAK.22 Hartree-Fock and global hybrid r2SCANX77 calculations were also performed, with percentages of exact Hartree-Fock exchange discussed in the text of the paper. Ground-state magnetic orderings reported experimentally (see Table 1), often derived from neutron diffraction experiments, were imposed in all simulations of MiOjs.
Geometries (coordinates, volumes, and cell shapes) converged when the forces on all atoms were lower than 0.01 eV/Å. All properties were calculated with r2SCAN geometries unless explicitly mentioned. For structure relaxation, a $\Gamma$-centered Monkhorst-Pack180 grid with a density of 48 k-points per Å for all systems. All other calculations used a k-grid of approximately 700 k-points per Å3.
@@
P.C. and J.P.P. designed and supervised the project. H.R.G. prepared and performed the simulations, data collection, and data analysis. P.C., J.P.P., J.S., R.Z., and Y.W. contributed to the data analysis. H.R.G. and P.C. wrote the first draft of the manuscript. All authors contributed to the discussion and the final version of this manuscript.
@@
The authors declare no competing interests.
@@ Acknowledgements.
The Welch Foundation is acknowledged for providing P.C. a Robert A. Welch Professorship at the Texas Center for Superconductivity and the Welch Foundation grant No. E-2227- 20250403. We are grateful for the support of the Research Computing Data Core at the University of Houston for assistance with the calculations carried out in this work. J.P.P. acknowledges support from the National Science Foundation under grant DMR-2426775 and the Department of Energy Office of Science under grant DE-SC001833. J.S. acknowledges support from the National Science Foundation under grant DMR-2042618.

### v3

@@ IV Conclusions
Despite its excellent performance in other areas,22, 153 the LAK meta-GGA has been found less accurate than r2SCAN for the oxidation energies (Fig. 2) and somewhere between r2SCAN and r2SCAN10 for the band gaps (Fig. 6) of these transition-metal oxides.
Note that the first, second, and third ionization energies of the 3d transition metal atoms also improve from LSDA to PBE to r2SCAN.173 That improvement is dramatic for the second and third ionization energies, which seem free from strong correlation effects.
The present investigation and other recent works82, 83, 84, 91, 93, 173 suggest that, in many but not all situations with main-group and transition-metal elements, the functional-driven errors of r2SCAN energy differences can far exceed the relatively small density-driven errors. This has the important consequence that significant improvements in the r2SCAN energy differences can be achieved by doing the expensive self-consistent iteration and geometry optimization with the computationally efficient r2SCAN meta-GGA (possibly including an efficient van der Waals dispersion corrections), and then applying a more expensive nonlocal functional (e.g., a hybrid or self-interaction correction) in a single post-self-consistent calculation, as we have done here in r2SCAN10@r2SCAN. All properties studied here improve on average from LSDA to PBE to r2SCAN to r2SCAN10@r2SCAN. Replacing a small fraction of r2SCAN exchange by the same fraction of exact Hartree-Fock exchange should have little or no effect on the satisfaction of exact constraints.
@@ V First-Principles Calculation Details
All calculations presented in this work are performed using the DFT formalism, as implemented in the Vienna Ab initio Simulation Package (VASP).174, 175, 176 The PAW potentials describe the core electrons.177, 178 The electrons from 3s, 3p, 3d, and 4s orbitals are explicitly considered for the transition metal atoms. Using a PAW set that treats fewer valence electrons explicitly (3p, 3d, and 4s) significantly impacts MiOj oxidation energies and band gaps (Supplementary Fig. 8). The kinetic energy cutoff for the plane waves was set to 700 eV, and the total energy was converged to 10-6 eV per cell. Various DFT exchange and correlation functionals were used in their collinear spin-polarized implementation. These are the LSDA-PW92,17 the Perdew, Burke, and Ernzerhof (PBE),18 the meta-GGA functionals r2SCAN,99 and LAK.22 Hartree-Fock and global hybrid r2SCANX77 calculations were also performed, with percentages of exact Hartree-Fock exchange discussed in the text of the paper. Ground-state magnetic orderings reported experimentally (see Table 1), often derived from neutron diffraction experiments, were imposed in all simulations of MiOjs.
Geometries (coordinates, volumes, and cell shapes) converged when the forces on all atoms were lower than 0.01 eV/Å. All properties were calculated with r2SCAN geometries unless explicitly mentioned. For structure relaxation, a $\Gamma$-centered Monkhorst-Pack179 grid with a density of 48 k-points per Å for all systems. All other calculations used a k-grid of approximately 700 k-points per Å3.
@@
P.C. and J.P.P. designed and supervised the project. H.R.G. prepared and performed the simulations, data collection, and data analysis. P.C., J.P.P., J.S., R.Z., Y.W., A.R., and A.P. contributed to the data analysis. H.R.G. and P.C. wrote the first draft of the manuscript. All authors contributed to the discussion and the final version of this manuscript.
@@
The authors declare that they have no competing interests.
@@ Acknowledgements.
The Welch Foundation is acknowledged for providing P.C. a Robert A. Welch Professorship at the Texas Center for Superconductivity and the Welch Foundation grant No. E-2227- 20250403. We are grateful for the support of the Research Computing Data Core at the University of Houston for assistance with the calculations carried out in this work. J.P.P. acknowledges support from the National Science Foundation under grant DMR–2426275 and the Department of Energy Office of Science under grant DE–SC–0018331. J.S. acknowledges support from the National Science Foundation under grant DMR–2042618.